OpenStudy (kainui):

What's the smallest number expressible as the sum of two cubes in two different ways?

OpenStudy (mathmath333):

what u mean by two diffferent ways

OpenStudy (kainui):

What is the smallest positive integer n to satisfy this for a,b,x,y all different.\[\Large n =a^3+b^3 = x^3+y^3\]

OpenStudy (kainui):

For instance 9 =1^3+2^3, but there's no other way to write 9 as the sum of two cubes, only this one way. =)

OpenStudy (mathmath333):

and a,b , x and y are all positive ?

OpenStudy (kainui):

I believe so.

OpenStudy (mathmath333):

\(1729\)

OpenStudy (kainui):

Show your work =)

OpenStudy (mathmath333):

http://en.wikipedia.org/wiki/1729_(number)

OpenStudy (mathmath333):

my work is too lazy =)

OpenStudy (kainui):

Meh, waste of time